OpenMath XML Translation Server Manual

Edition : auto generated by oxgentexi on 23 February 2009

OpenXM.org

1 OpenMath Functions(Version 1999)

The functions in this section is defined in the file ‘om.rr’. An environment to execute Java codes must be set to call the functions described in this section. Author of OMproxy : Yasushi Tamura.

1.0.1 om_start
om_start()

:: Start OMproxy server to make a translation between CMO and OpenMath XML (CD’s in 1999) expressions.

return Number

[155] load("om.rr");
1
[160] om_start();
control: wait OX
Trying to connect to the server... Done.
0
[161] om_xml(<<1,0>>+2*<<0,1>>);
<OMOBJ><OMA><OMS name="DMP" cd="poly"/>
<OMA><OMS name="PolyRing" cd="poly"/>
 <OMI>2</OMI></OMA><OMA>
 <OMS name="SDMP" cd="poly"/>
 <OMA><OMS name="Monom" cd="poly"/><OMI>1</OMI><OMI>1</OMI><OMI>0</OMI></OMA>
 <OMA><OMS name="Monom" cd="poly"/><OMI>2</OMI><OMI>0</OMI><OMI>1</OMI></OMA>
</OMA></OMA></OMOBJ>
[162] om_xml_to_cmo(@);
(1)*<<1,0>>+(2)*<<0,1>>
1.0.2 om_xml
om_xml(s|proc=p)

:: Translate CMO expression of s to a XML expression of OpenMath(CD’s in 1999).

return String

p Number

s Object

• Translate CMO s to a XML expression of OpenMath(CD’s in 1999).

For (I=0; I<10; I++) {
   A = 2^I;
   B = om_xml(A);
   C = om_xml_to_cmo(B);
   print(A == C);
}

1.0.3 om_xml_to_cmo
om_xml_to_cmo(s|proc=p)

:: Translate XML expression (CD’s in 1999) s of OpenMath to a CMO.

return Object p Number s String

• Translate XML expression (CD’s in 1999) s of OpenMath to a CMO.

Index 3

Index

(Index is nonexistent) (Index is nonexistent)

Short Contents

1 OpenMath Functions(Version 1999) . . . . . . . . . . . . . . . . . . . 1 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Table of Contents

1 OpenMath Functions(Version 1999) . . . . . . . . . 1

1.0.1 om_start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.0.2 om_xml . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.0.3 om_xml_to_cmo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3